3. Among the following which can not be an electric field? E~ = c(xyi +2yzj +3xzk), E~ = c(y^2i + (2xy + z^2)j + 2yzk), where c is a constant.

What would be the dimension of c? (Problem 2.20 of Griffiths, 4th

edition)


1
Expert's answer
2021-09-18T14:58:51-0400

Given:

"{\bf E}_1 = c(xy{\hat i} +2yz{\hat j} +3xz{\hat k})"

"{\bf E}_2 = c(y^2{\hat i} +(2xy+z^2){\hat j} +2yz{\hat k})"


The electrostatic field must satisfy the condition

"\rm rot\:{\bf E}=\bf 0"

Let's check it

"\rm rot\:{\bf E}_1=\begin{vmatrix}
 \hat i & \hat j &\hat k\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
E_x&E_y&E_z
\end{vmatrix}"

"=c\begin{vmatrix}
 \hat i & \hat j &\hat k\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
xy&2yz&3xz
\end{vmatrix}=c(-2y\hat i-3z\hat j-y\hat z)\neq0"

"{\rm rot\:{\bf E}_2}=c\begin{vmatrix}
 \hat i & \hat j &\hat k\\
 \frac{\partial}{\partial x} & \frac{\partial}{\partial y} & \frac{\partial}{\partial z} \\
y^2&2xy+z^2&2yz
\end{vmatrix}"

"=c(0\hat i+0\hat j+0\hat z)=\bf 0"

Answer: "E_1" can not be an electric field.

The dimention of constant "c" is "\rm V/m^3"